OpenStudy (anonymous):

A fireworks company has two types of rockets called Zinger 1 and Zinger 2. The polynomial -16t^2+150t gives the height in feet of Zinger 1 at t seconds after launch. The polynomial: -16t^2 + 165t gives the height of Zinger 2 at t seconds after launch. If the rockets are launched at the same time and both explode 6 seconds after launch, how much higher is Zinger 2 than Zinger 1 when they explode?

OpenStudy (anonymous):

would this be the right formula ? t=6 Zinger 1: -16 * t^2 + 150 * t = -16 * 6^2 + 150 * 6 =324 Zinger 2: -16 * t^2 + 165 * t = -16 * 6^2 + 165 * 6 = 414 414-324=90

OpenStudy (campbell_st):

your solution is correct

OpenStudy (anonymous):

thank you